A train is moving along a track with constant speed v1 relative to the ground. A person on the train holds a ball of mass m and throws it toward the front of the train with a speed v2 relative to the train.

Calculate the change in kinetic energy of the ball in the Earth frame of reference. Express your answer in terms of the variables m, v1, and v2.

Calculate the change in kinetic energy of the ball in the train frame of reference. Express your answer in terms of the variables m, v1, and v2.

How much work was done on the ball in the Earth frame of reference? Express your answer in terms of the variables m, v1, and v2.

How much work was done on the ball in the train frame of reference?Express your answer in terms of the variables m, v1, and v2.
Answer #1

velocity of ball in train reference = v2

velocity of ball in earth reference = v1+v2

PART(A).

ke1 = 1/2mv12

ke2 =1/2m (v1+v2)2

change in KE =ke2-ke1

=1/2m(v22+2v1v2)

Part (B).

train frame of reference

ke1=0

Ke2=1/2mv22

Difference in KE = 1/2mv22

part (C).

work done = change in kinectic enrgy =1/2m(V22+2v1v2)

part (D).

work done = change in kinectic enrgy=1/2mv22